# Latent Diffeomorphic Dynamic Mode Decomposition

### v1

@@ 1 Introduction
In scenarios without memory where a one-to-one mapping exists, Dynamic Mode Decomposition (DMD) schmid2010dynamic has proven highly effective as an interpretable data reduction method. Its success has led to numerous extensions tu2014on ; dawson2016characterizing ; hemati2017biasing . Recent advancements have focused on adapting DMD to non-linear cases, primarily through Koopman operator theory. This approach typically follows one of two paths: using non-linear dictionaries williams2015data ; li2017extended or employing non-linear mappings takeishi2017learning ; lusch2018deep ; meng2024koopman ; hou2024invertible , with the latter being more popular. A key challenge in the non-linear mapping approach is the need for an (approximate) inverse function to recover the actual state from the dynamics in the mapped domain. While early attempts used approximations takeishi2017learning ; lusch2018deep , more recent research has shown that diffeomorphisms meng2024koopman ; hou2024invertible offer improved results.
@@ 2 General Diffeomorphic Dynamic Mode Decomposition
We note that this approach can be seen as a combination of the approaches taken in lusch2018deep – where a similar parametrization for $\mathbf{K}$ is used, but $\varphi^{-1}$ and $\varphi$ are replaced by (non-invertible) neural networks – and state-of-the-art approaches taken in meng2024koopman ; hou2024invertible – where diffeomorphisms are used, but no block diagonal assumptions are made on $\mathbf{K}$.
@@ 3 Latent Diffeomorphic Dynamic Mode Decomposition
where $\mathbf{K}_{1}:\mathbb{R}^{m\times m}$ and $\mathbf{K}_{2}:\mathbb{R}^{d\times d}$ are block-diagonal matrices with blocks of the form (3) [^1] Note that for this to be well-defined, we need both $m$ and $d$ to be even., $g:\mathbb{R}^{d}\to\mathbb{R}^{n}$ is a regression mapping, and the latent variable is initialized by some $\mathbf{p}^{0}\in\mathbb{R}^{d}$.
@@
and aim to learn the regression mapping $g$, the diffeomorphism $\varphi$, the coupling $f$, and the initialization $\mathbf{q}^{0}$ along with the $\mu_{j},\omega_{j}$ that parametrize $\mathbf{K}$ through (3).
@@
The latent dynamics $(\mathbf{K})^{i}\mathbf{q}^{0}$ encode the memory of the system and is evolving independently of $\mathbf{x}^{i}$. In other words, we really have a non-one-to-one mapping from $\mathbf{x}^{i}$ to $\mathbf{v}^{i}$, where the memory is encoded in a more interpretable way than for RNN-like methods, with the key difference that we explicitly have to solve for an initial condition $\mathbf{q}_{0}$. Having said that, standard RNNs famously have trouble with maintaining information for longer terms and can be unstable. Both of these issues can be attributed to vanishing or exploding gradients, which LDDMD can also suffer from for $\mu\neq 0$. So in practice, we assume for LDDMD that $\mu=0$.
@@ 4 Training
In this case, we found the following parameterization suitable: (i) the diffeomorphism as an invertible neural network composed of a single additive coupling layer dinh2014nice with learnable polynomial activation on either the even or odd indices, initialized such that $\varphi_{\theta^{(0)}}(\mathbf{x})=\mathbf{x}$; (ii) the coupling as a feedforward network with learnable polynomial activation functions, initialized such that $f_{\eta^{(0)}}(\mathbf{x})\approx\mathbf{0}$; (iii) the frequencies $\omega_{j}$ as learnable parameters, initialized as the $d/2$ most prevalent frequencies in the Fourier spectrum [^2] Factoring in the time $\Delta t$ between observations. of the 1D signal $\{\mathbf{v}^{i}\}_{i=0}^{N}$; (iv) the latent initialization as a learnable vector, initialized as $\mathbf{q}^{(0)}_{0}=\mathbf{0}$; (v) and the regression $g_{\xi}$ as a feedforward network with softplus activation functions and random initialization.
@@ 6 Conclusions
In conclusion, this work presents Latent Diffeomorphic Dynamic Mode Decomposition (LDDMD), which offers a novel approach to data reduction and prediction of complex time series by combining the strengths of Dynamic Mode Decomposition (DMD) and Recurrent Neural Networks (RNNs). LDDMD effectively handles non-linear systems with memory, providing both interpretability and predictive power, as demonstrated through successful streamflow prediction experiments. However, our results suggest that further adaptations are needed for state-of-the-art performance. Future work will focus on improving LDDMD’s robustness to noise and evaluating its ability to capture rare events compared to LSTM-based approaches.
@@ Appendix B Training details
    • Regression Mappings: Multi layer perceptron with one hidden layer with dimension $m_{g}$ and softplus activation.

### v2

@@ 1 Introduction
In scenarios without memory where a one-to-one mapping exists, Dynamic Mode Decomposition (DMD) schmid2010dynamic has proven highly effective as an interpretable data reduction method. Its success has led to numerous extensions tu2014on ; dawson2016characterizing ; hemati2017biasing . Recent advancements have focused on adapting DMD to non-linear cases, primarily through Koopman operator theory Mezic2021 . This approach typically follows one of two paths: using non-linear dictionaries williams2015data ; li2017extended or employing non-linear mappings takeishi2017learning ; lusch2018deep ; meng2024koopman ; hou2024invertible , with the latter being more popular. A key challenge in the non-linear mapping approach is the need for an (approximate) inverse function to recover the actual state from the dynamics in the mapped domain. While early attempts used approximations takeishi2017learning ; lusch2018deep , more recent research has shown that diffeomorphisms meng2024koopman ; hou2024invertible offer improved results.
@@ 2 General Diffeomorphic Dynamic Mode Decomposition
We note that this approach can be seen as: (i) a combination of the approaches taken in lusch2018deep – where a similar parametrization for $\mathbf{K}$ is used, but $\varphi^{-1}$ and $\varphi$ are replaced by (non-invertible) neural networks – and state-of-the-art approaches taken in meng2024koopman ; hou2024invertible – where diffeomorphisms are used, but no block diagonal assumptions are made on $\mathbf{K}$, and (ii) a special case of Bollt2018 where the eigenvectors of the Koopman operator are only assumed to generate a local diffeomorphism.
@@ 3 Latent Diffeomorphic Dynamic Mode Decomposition
where $\mathbf{K}_{1}:\mathbb{R}^{m\times m}$ and $\mathbf{K}_{2}:\mathbb{R}^{d\times d}$ are block-diagonal matrices with blocks of the form (3) [^1] Note that for this to be well-defined, we need both $m$ and $d$ to be even., $g:\mathbb{R}^{d}\to\mathbb{R}^{n}$ is a mapping, and the latent variable $\mathbf{p}$ is initialized by some $\mathbf{p}^{0}\in\mathbb{R}^{d}$.
@@
and aim to learn the mapping $g$, the diffeomorphism $\varphi$, the coupling $f$, and the initialization $\mathbf{q}^{0}$ along with the $\mu_{j},\omega_{j}$ that parametrize $\mathbf{K}$ through (3).
@@
The latent dynamics $(\mathbf{K})^{i}\mathbf{q}^{0}$ encode the memory of the system and is evolving independently of $\mathbf{x}^{i}$. In other words, we really have a non-one-to-one mapping from $\mathbf{x}^{i}$ to $\mathbf{v}^{i}$, where the memory is encoded in a more interpretable way than for RNN-like methods, with the key difference that we explicitly have to solve for an initial condition $\mathbf{q}_{0}$. Having said that, standard RNNs famously have trouble with maintaining information for longer terms and can be unstable. Both of these issues can be attributed to vanishing or exploding gradients, which LDDMD can also suffer from for $\mu_{j}\neq 0$ in $\mathbf{K}$, see (3). So in practice, we assume for LDDMD that $\mu_{j}=0$.
@@ 4 Training
In this case, we found the following parameterization suitable: (i) the diffeomorphism as an invertible neural network composed of a single additive coupling layer dinh2014nice with learnable polynomial activation on either the even or odd indices, initialized such that $\varphi_{\theta^{(0)}}(\mathbf{x})=\mathbf{x}$; (ii) the coupling as a feedforward network with learnable polynomial activation functions, initialized such that $f_{\eta^{(0)}}(\mathbf{x})\approx\mathbf{0}$; (iii) the frequencies $\omega_{j}$ as learnable parameters, initialized as the $d/2$ most prevalent frequencies in the Fourier spectrum [^2] Factoring in the time $\Delta t$ between observations. of the 1D signal $\{\mathbf{v}^{i}\}_{i=0}^{N}$; (iv) the latent initialization as a learnable vector, initialized as $\mathbf{q}^{(0)}_{0}=\mathbf{0}$; (v) and the mapping $g_{\xi}$ as a feedforward network with softplus activation functions and random initialization. Limiting the diffeomorphism to only use additive coupling layers serves as a form of regularization, helping to prevent the mapping from producing excessively large or small output values.
@@ 6 Conclusions
In conclusion, this work presents Latent Diffeomorphic Dynamic Mode Decomposition (LDDMD), a novel method that combines Dynamic Mode Decomposition and Recurrent Neural Networks for data reduction and nonlinear time series prediction. LDDMD offers both interpretability and strong performance, as shown in streamflow forecasting, though further improvements are needed for state-of-the-art results.
@@ Appendix B Training details
Based on our experience, both shallow and deep networks delivered strong performance, provided the activation functions were regular. For instance, non-regular ReLU activation performed poorly. We prioritized simplicity in our models, which primarily guided the selection of the parameters reported below.
@@
    • Mappings: Multi layer perceptron with one hidden layer with dimension $m_{g}$ and softplus activation.